OpenStudy (anonymous):

Find an equation of the line through (2,5) and parallel to y=4x-4. write the equation using the function notation.

OpenStudy (ajprincess):

Let f(x)=y, where y=mx+c be the equation of the line. m-slope c-y_intercept For parallel lines slopes are equal. Slope of the given line=4. So slope m of the required line=4 y=4x+c Nw plug in 2 in place of x and 5 in place of y to find c. Can u find it @NeedSomehelpplz?

OpenStudy (anonymous):

y=8+5 ?

OpenStudy (anonymous):

y=8x+5 ?

OpenStudy (ajprincess):

hmm no. Let us find c first by substituting the values of x nd y in y=4x+c 5=4*2+c c=8-5 =3 Is this clear @Needsomehelpplz?

OpenStudy (anonymous):

Thanks, I think I've got it.

OpenStudy (ajprincess):

Nw ur equation will be y=4x+3 In function form f(x)=4x+3 Welcome:)

OpenStudy (ajprincess):

If something is not clear plz ask:)